# X-Boundary: Establishing Exact Safety Boundary to Shield LLMs from Multi-Turn Jailbreaks without Compromising Usability

### v2

@@ Abstract
Despite the rapid development of safety alignment techniques for LLMs, defending against multi-turn jailbreaks is still a challenging task. In this paper, we conduct a comprehensive comparison, revealing that some existing defense methods can improve the robustness of LLMs against multi-turn jailbreaks but compromise usability, i.e., reducing general capabilities or causing the over-refusal problem. From the perspective of mechanism interpretability of LLMs, we discover that these methods fail to establish a boundary that exactly distinguishes safe and harmful feature representations. Therefore, boundary-safe representations close to harmful representations are inevitably disrupted, leading to a decline in usability. To address this issue, we propose X-Boundary to push harmful representations away from boundary-safe representations and obtain an exact distinction boundary. In this way, harmful representations can be precisely erased without disrupting safe ones. Experimental results show that X-Boundary achieves state-of-the-art defense performance against multi-turn jailbreaks, while reducing the over-refusal rate by about 20% and maintaining nearly complete general capability. Furthermore, we theoretically prove and empirically verify that X-Boundary can accelerate the convergence process during training. Please see our code at: https://github.com/AI45Lab/X-Boundary. Warning: this paper includes examples that may be offensive or harmful.
Xiaoya Lu1,2,*Dongrui Liu1,*Yi Yu2Luxin Xu3Jing Shao2,†
1 School of Electronic Information and Electric Engineering, Shanghai Jiao Tong University
2 Shanghai Artificial Intelligence Laboratory
3 University of Electronic Science and Technology of China
footnotetext: Correspondence to: Jing Shao <shaojing@pjlab.org.cn>
@@ 1 Introduction
As large language models (LLMs) have demonstrated impressive abilities (OpenAI, 2024; AI@Meta, 2024; Team, 2023; Chen et al., 2024) and are increasingly deployed in diverse real-world applications (Zhou et al., 2024a; Huang et al., 2024), their security vulnerabilities have raised growing concern. One of the most significant security threats is “jailbreaks”, where deliberately designed prompts are used to elicit harmful responses from LLMs (Yi et al., 2024; Liu et al., 2023). Although current safety alignment techniques (Rafailov et al., 2024; Yuan et al., 2024) can defend against various single-turn jailbreaks, multi-turn jailbreaks remain a challenge. Unlike single-turn attacks, multi-turn jailbreaks exploit flexible multi-turn dialogues to bypass the safeguards of LLMs (Zhou et al., 2024b; Liu et al., 2024b), making them difficult to detect and defend against.
In this paper, we pioneeringly adapt and comprehensively compare four single-turn defense methods against multi-turn jailbreaks, including supervised fine-tuning (SFT) (Ren et al., 2024b), Direct Preference Optimization (DPO) (Jiang et al., 2024), Gradient Ascent (GA) (Zhang et al., 2024b), and circuit breaking (CB) (Zou et al., 2024). Although these methods can reduce the attack success rate (ASR) of multi-turn jailbreaks, they compromise the usability of LLMs, i.e., reducing general capability or leading to severe over-refusal issues. For example, the coding capability relatively reduces about 20% after SFT and the over-refusal rate exceeds 50% after GA. None of them successfully achieve a balance between defense robustness and usability. Over-refusal means that LLMs may unreasonably reject safe prompts (Röttger et al., 2023; Shi et al., 2024; Cui et al., 2024). For instance, Fig. 1 shows that LLMs refuse to answer a safe prompt “What are the harmful effects of meth on the human body?”.
From the perspective of the internal mechanism of LLMs, establishing a boundary to exactly distinguish harmful and safe feature representations is crucial for achieving the balance between robustness and usability. This distinction boundary enables harmful representations to be intervened in while safe representations are preserved. However, existing multi-turn defense methods fail to learn an exact boundary due to the lack of an explicit boundary formulation, as shown in Fig. 1. Some boundary-safe queries, e.g., “What are the harmful effects of meth on the human body?” involve sensitive information, e.g., “meth”, and their representations are close to harmful representations. In this way, these boundary-safe queries are mistakenly treated as harmful and rejected by LLMs.
@@
To strike a balance between defense robustness and usability, we propose an X-Boundary that explicitly formulates the boundary between harmful and safe representations. Specifically, X-Boundary optimizes the LLM to push harmful representations far away from boundary-safe representations, while keeping trained boundary-safe representations close to their original representations. In this way, we obtain a precise distinction boundary and these harmful representations are further erased. Experimental results demonstrate that X-Boundary decreases the ASR of multi-turn jailbreaks from 58.5% to 16.5% on Llama-3-8B-Instruct, while reducing the over-refusal rate by an average of 20.5% compared to state-of-the-art (SOTA) method and preserving nearly complete general capability. Additionally, we theoretically analyze the feature learning trend of LLM with X-Boundary from the perspective of optimal transport theory. Theoretical analysis and experimental results indicate that X-Boundary achieves about 22% improvement in the learning speed.
X-Boundary achieves a win-win outcome with enhanced robustness against multi-turn jailbreaks and minimal decline in usability. We do not expect X-Boundary can independently address all security threats without collaborating with alignment methods such as SFT and Reinforcement Learning from Human Feedback (RLHF). Instead, we believe that X-Boundary can provide a more efficient and fine-grained defense for safety-aligned LLMs, ultimately enhancing the prospects of deploying robust AI systems in diverse real-world applications.
Table 1: Comparison of existing defense methods and X-Boundary.
Models Methods Multi-Turn ASR (%) $\downarrow$ Over-Refusal Rate (%) $\downarrow$ General Capability (%) $\uparrow$
Llama-3- 8B-Instruct Vanilla 58.50 25.00 34.00 6.80 9.00 8.00 13.67 68.30 79.08 59.18
SFT 19.50 0.50 8.00 27.20 42.33 22.00 57.33 68.17 76.19 54.27
DPO 17.50 5.00 14.00 20.00 28.33 17.33 41.00 68.01 75.59 58.54
GA 38.50 1.50 12.00 10.80 15.00 13.33 35.33 68.25 77.86 62.20
CB 16.50 0.50 10.00 23.60 27.67 36.00 52.00 67.66 78.47 59.76
X-Boundary 16.50 1.00 10.00 8.40 14.00 8.00 28.67 67.94 78.70 59.76
Qwen2.5 -7B-Chat Vanilla 76.00 39.50 62.00 6.00 19.33 1.67 5.60 74.26 80.67 81.71
SFT 21.00 6.00 18.00 46.00 57.67 29.33 53.67 74.30 76.42 77.44
DPO 38.00 12.00 24.00 21.60 25.67 11.67 32.33 73.63 80.97 80.49
GA 38.00 21.00 12.00 58.40 70.00 67.67 85.33 74.58 80.43 79.27
CB 15.50 5.50 12.00 20.60 26.00 34.00 43.67 74.21 80.36 81.10
X-Boundary 17.50 7.50 16.00 10.40 16.67 5.33 15.00 74.17 80.52 81.10
Mistral-7B -Instruct-v0.2 Vanilla 70.00 49.50 40.00 10.00 21.00 4.33 13.00 59.98 45.34 34.76
SFT 37.50 22.00 18.00 53.60 42.00 29.33 58.67 58.94 41.55 27.44
DPO 44.50 19.00 28.00 25.20 38.67 20.33 37.67 58.79 43.21 34.76
GA 24.00 9.00 10.00 38.40 50.67 35.67 71.33 60.13 45.64 34.76
CB 15.00 11.50 12.00 45.20 32.33 55.00 50.00 59.91 46.63 33.54
X-Boundary 16.00 13.50 14.00 19.20 23.33 10.34 26.33 59.83 45.34 36.59
2 Comparison of Existing Defense Methods Against Multi-Turn Jailbreaks
To the best of our knowledge, we are the first to conduct a comprehensive evaluation of classic defense approaches against multi-turn jailbreaks, considering both defense robustness and impact on usability. Although previous studies (Ren et al., 2024b; Jiang et al., 2024) have employed Supervised Fine-Tuning (SFT) and Direct Preference Optimization (DPO) in multi-turn defense scenarios, many other defense methods effective against single-turn jailbreaks, such as Gradient Ascent (GA) and Circuit Breaking (CB), have been overlooked and have not been thoroughly compared. In Section 2.1, we outline the process of constructing training data, reproducing SFT and DPO, and adapting GA and CB for multi-turn defense scenarios. In Section 2.2, we present and analyze the evaluation results, revealing the shortcomings of existing defense methods in balancing robustness and usability.
2.1 Adaption and Evaluation of Single-Turn Defense Methods Against Multi-Turn Jailbreaks
We compare the following defense methods against multi-turn jailbreak (Ren et al., 2024b; Jiang et al., 2024) on Llama-3-8B-Instruct (AI@Meta, 2024), Qwen2.5-7B-Chat (Yang et al., 2024a), and Mistral-7B-Instruct-v0.2 (Jiang et al., 2023):
  • SFT (Ren et al., 2024b): fine-tuning LLMs using harmful queries as inputs and refusal answers as supervised labels directly.
  • DPO (Rafailov et al., 2024; Jiang et al., 2024): aligning LLMs using harmful queries as inputs, harmful answers as rejected responses, and refusal answers as chosen responses.
  • GA (Zhang et al., 2024b; Lu et al., 2024a): unlearning harmful knowledge by training with gradient ascent optimization methods .
  • CB (Zou et al., 2024): remapping the representations of harmful knowledge to desired targeted representations.
Construct multi-turn defense datasets. We construct the multi-turn defense training datasets based on SafeMTData (Ren et al., 2024b). SafeMTData consists of 1680 safe multi-turn dialogues for the safety alignment of LLMs in multi-turn interactions. For SFT, we directly exploit SafeMTData as a multi-turn training dataset following Ren et al. (2024b). For DPO, we curate harmful responses to the harmful multi-turn queries in SafeMTData and constructed a multi-turn preference dataset following Jiang et al. (2024). For CB and GA, to remove harmful knowledge that could be elicited through multi-turn attacks, we add pairs of harmful queries from SafeMTData along with the curated harmful responses into their respective defense training datasets. (Zhang et al., 2024b; Zou et al., 2024) More details about data construction are illustrated in Appendix B.1 and training settings of these methods are listed in Appendix B.2.
Evaluation of defense robustness. To evaluate the robustness of these defense methods, we test them against three SOTA multi-turn jailbreak attacks: ActorAttack (Ren et al., 2024b), RedQueen (Jiang et al., 2024) and Crescendo (Russinovich et al., 2024). The evaluation metric used is the Attack Success Rate (ASR), defined as the proportion of attack attempts that successfully elicit harmful content from the LLMs. A lower ASR indicates greater robustness against multi-turn jailbreak attacks. We also evaluate the compliance rates of LLMs with harmful requests in HarmBench and the results are listed in the Appendix A.1.
Evaluation of usability. We evaluate the impact of defense methods on the usability of LLMs from two perspectives: over-refusal and the decline of general capability. Over-refusal is evaluated using XSTest (Röttger et al., 2023), OKTest (Shi et al., 2024), OR-Bench (Cui et al., 2024), and PHTest (An et al., 2024). The corresponding evaluation metric is the over-refusal rate, which measures the proportion of benign prompts that the model incorrectly refuses to answer. A lower over-refusal rate indicates better usability. General capability, including general knowledge, mathematics, and coding ability, is evaluated using MMLU (Hendrycks et al., 2020), GSM8K (Cobbe et al., 2021) and HumanEval (Chen et al., 2021), respectively. Please see Appendix B.4 for more details about evaluation.
2.2 Experimental Results and Analysis
Existing methods fail to strike a balance between robustness and usability. Table 1 shows that existing methods can effectively reduce the ASR of multi-turn jailbreaks after training with the aforementioned data. However, SFT, DPO, and GA even tend to severely compromise general capabilities when achieving good performance, commonly referred to as the “alignment tax” (Ouyang et al., 2022b). For instance, SFT results in a 3%-7% decrease in both the mathematical and the coding abilities of the three models. Moreover, all of these methods lead to a significant increase in the over-refusal rate.
@@ 2.2 Experimental Results and Analysis
An excessively high over-refusal rate makes the ASR unreliable. In PHtest, the over-refusal rate of Llama-3-8B increases from 13.67% to more than 40% on all methods. The over-refusal rates after GA are more than 50% on four benchmarks. The high over-refusal rate reflects that these methods cannot precisely distinguish harmful queries and build effective defense mechanisms for them. Instead, they simply reduce the ASR of multi-turn attacks by indiscriminately rejecting input queries, which is not trustworthy and undermines the model’s usability in real-world scenarios. Therefore, it is necessary to analyze the cause of over-refusal and propose a more precise defense method to mitigate it while preserving robustness against multi-turn jailbreak.
3 X-Boundary: Optimize Exact Boundary to Balance Defense Robustness and Usability
@@
In this section, we propose X-Boundary to balance robustness against multi-turn jailbreaks and usability by explicitly formulating the distinction boundary. Section 3.1 analyzes the essential mechanism of decline in usability. Section 3.2 introduces the optimization objective of X-Boundary. Section 3.3 theoretically proves that X-Boundary may ease the learning difficulty and contribute to fast learning.
@@
Analysis of safety-usability trade-off from the perspective of interpretability mechanism. Existing defense methods (Zou et al., 2024, 2023) typically improve the adversarial robustness of LLMs by intervening harmful feature representations $\{\mathcal{R}_{\mathcal{M}}\left(x_{i}^{h}\right)\}_{i=1}^{N}$. Specifically, SFT (Yuan et al., 2024) and CB (Zou et al., 2024) remap harmful representations to refusal representations $\mathcal{R}_{\mathcal{M}}\left(x_{i}^{r}\right)$. In this process, these methods implicitly train LLMs to learn a boundary that distinguishes harmful representations and safe representations $\{\mathcal{R}_{\mathcal{M}}\left(x_{i}^{s}\right)\}_{i=1}^{N}$. However, Fig. 2 shows that the boundary learned through this implicit training is imprecise, with some boundary-safe representations $\{\mathcal{R}_{\mathcal{M}}\left(x_{i}^{b}\right)\}_{i=1}^{N}$ mixed with harmful representations rather than being clearly distinguished. In this way, these boundary-safe representations are mistakenly treated as harmful ones, leading LLMs to refuse the corresponding boundary-safe queries and ultimately reducing usability.
@@
3.2 Explicit Formulation for Distinction Boundary
We propose X-Boundary to explicitly formulate the distinction boundary between safe and harmful representations. The key idea is to push harmful representations far away from boundary-safe representations through explicit loss function, such that harmful representations can be effectively and precisely erased without disrupting safe ones. In this way, a balance between defense robustness and LLM usability can be achieved.
@@ 3.2 Explicit Formulation for Distinction Boundary
To explicit formulate a precise distinction boundary, we propose separate loss $\mathcal{L}_{\texttt{s}}$ to increase the distance between harmful representations $\{\mathcal{R}_{\mathcal{M}_{\theta}}\left(x_{i}^{h}\right)\}_{i=1}^{N}$ and boundary-safe representations $\{\mathcal{R}_{\mathcal{M}_{\theta}}\left(x_{i}^{b}\right)\}_{i=1}^{N}$. Since most $\{\mathcal{R}_{\mathcal{M}_{\theta}}\left(x_{i}^{h}\right)\}_{i=1}^{N}$ will be remapped to $\{\mathcal{R}_{\mathcal{M}_{\theta}}\left(x_{i}^{r}\right)\}_{i=1}^{N}$ due to the following erasure operation, we can separate them by directly optimizing $\mathcal{R}_{\mathcal{M}_{\theta}}\left(x_{i}^{r}\right)$ to be orthogonal to $\mathcal{R}_{\mathcal{M}_{\theta}}\left(x_{i}^{b}\right)$ as shown in Fig. 3:
$\mathcal{L}_{\texttt{s}}=\frac{1}{\left|S_{s}\right|}\sum_{i=1}^{\left|S_{s}\right|}\texttt{ReLU}\left(\texttt{cos}\left(\mathcal{R}_{\mathcal{M}_{\theta}}\left(x_{i}^{r}\right),\mathcal{R}_{\mathcal{M}}\left(x_{i}^{b}\right)\right)\right)$ (1)
To establish robust defense against multi-turn attacks, we utilize erase loss $L_{e}$ to erase the representations of harmful QA pairs in $D_{\texttt{e}}$. $L_{e}$ optimizes $\mathcal{R}_{\mathcal{M}_{\theta}}\left(x_{i}^{h}\right)$ to be orthogonal to their original representations $\mathcal{R}_{\mathcal{M}_{\texttt{ref}}}\left(x_{i}^{h}\right)$ following (Zou et al., 2024):
@@
where $x_{i}$ represents a sample in retain set ($x_{i}\in D_{r}$), and $\mathcal{R}_{\mathcal{M}_{\theta}}$ and $\mathcal{M}_{\texttt{ref}}$ denote the model under training and the reference model before training, respectively.
To preserve usability of LLMs, we use retain loss $\mathcal{L}_{\texttt{r}}$ to maintain safe representations of data points in $S_{\texttt{r}}$. $\mathcal{L}_{\texttt{r}}$ minimizes the $\ell_{2}$ distance between trained representations and their original representations:
$\mathcal{L}_{\text{r}}=\frac{1}{\left|S_{r}\right|}\sum_{i=1}^{\left|S_{r}\right|}\left\|\mathcal{R}_{\mathcal{M}_{\theta}}\left(x_{i}\right)-\mathcal{R}_{\mathcal{M}}\left(x_{i}\right)\right\|_{2}$ (3)
Notably, to maintain the existing refusal mechanism of LLMs, refusal responses $x_{r}$ to harmful queries are added into $S_{\texttt{r}}$. Therefore, most $\{\mathcal{R}_{\mathcal{M}}\left(x_{h}\right)\}_{i=1}^{N}$ are finally optimized to refusal representations $\{\mathcal{R}_{\mathcal{M}}\left(x_{r}\right)\}\}_{i=1}^{N}$ under the joint effect of $\mathcal{L}_{\texttt{e}}$ and $\mathcal{L}_{\texttt{r}}$.
@@
where $c_{r}$, $c_{e}$ and $c_{s}$ are adaptive loss coefficients following (Zou et al., 2024; Ocampo et al., 2024). With the above optimization objective, X-Boundary can perform fine-grained optimization in the representation space to strike a balance between defense robustness and the usability of LLMs.
Algorithm 1 The optimization process of X-Boundary
1:  for $t=1$ to $T$ do
3:     Sample $\{(x_{i}^{b},x_{i}^{r})\}_{i=1}^{n}\sim\mathcal{D}_{b}$
4:     $c_{r}=\alpha\frac{t}{\beta}$, $c_{e}=c_{s}=\alpha(1-\frac{t}{\beta})$
5:     $\mathcal{L}_{\text{r}}=\frac{1}{n}\sum_{i=1}^{n}\left\|\mathcal{R}_{\mathcal{M}_{\theta}}\left(x_{i}\right)-\mathcal{R}_{\mathcal{M}}\left(x_{i}\right)\right\|_{2}$
8:     $\mathcal{L}=c_{r}\mathcal{L}_{r}+c_{e}\mathcal{L}_{e}+c_{s}\mathcal{L}_{s}$
9:     Update parameters $\theta$ to minimize $\mathcal{L}$
10:  end for

### v3

# X-Boundary: Establishing Exact Safety Boundary to Shield LLMs from Jailbreak Attacks without Compromising Usability

@@ Abstract
With the widespread application of large language models (LLMs) across various domains, techniques for enhancing their security have progressed rapidly. In this paper, we reveal that although existing defense methods can improve the robustness of LLMs against jailbreaks, they compromise usability, i.e., reducing general capabilities or causing the over-refusal problem. From the perspective of LLM mechanism interpretability, we discover that these methods fail to establish a boundary that exactly distinguishes safe and harmful feature representations. Therefore, boundary-safe representations close to harmful representations are inevitably disrupted, leading to a decline in usability. To address this issue, we propose X-Boundary to push harmful representations away from boundary-safe representations and obtain an exact distinction boundary. In this way, harmful representations can be precisely erased without disrupting safe ones. Experimental results show that X-Boundary achieves state-of-the-art defense performance against both single-turn and multi-turn jailbreak attacks, while reducing the over-refusal rate by about 20% and maintaining nearly complete general capability. Furthermore, we theoretically prove and empirically verify that X-Boundary can accelerate the convergence process during training. Code are released under https://github.com/AI45Lab/X-Boundary.
Warning: this paper includes examples that may be offensive or harmful.
@@ 1 Introduction
As large language models (LLMs) are increasingly deployed across diverse real-world applications gpt4; internlm; agent_survey, concerns about their security vulnerabilities have grown. One of the most serious threats is “jailbreaks”, where deliberately crafted single-turn or multi-turn prompts are used to elicit harmful responses from LLMs chao2023PAIR; autodan. To address this issue, several defense methods have been proposed defense_survey; zeng2024autodefense to enhance the robustness of LLMs against jailbreak attacks. However, these methods compromise the usability of LLMs, i.e., reducing general capability or leading to serious over-refusal issues alignment_tax; panda2024llm; orbench. Over-refusal means that LLMs may unreasonably reject prompts that are safe xstest; oktest. For instance, Fig. 1 shows that LLMs refuse to answer the harmless prompt “What are the harmful effects of meth on the human body?” merely due to the presence of the sensitive term “meth.”
In this paper, we conduct a comprehensive comparison of four widely used defense methods and clearly illustrate the trade-off between defense robustness and LLM usability. For example, Supervised Fine-Tuning (SFT) and Gradient Ascent (GA) lead to an alarmingly high over-refusal rate (ORR), often exceeding 50%, along with an approximate 20% drop in coding performance. Moreover, we find that training with multi-turn defense data severely exacerbates the trade-off, e.g., the ORR increases from 15% to 44%, while the ASR decreases from 30% to 12%. Although prior works have shown that certain methods oktest; wang2024surgical can alleviate over-refusal, our results show that these approaches weaken defense robustness, failing to resolve the trade-off.
Inspired by representation engineering zou2023representation, we investigate the root cause of the trade-off from the perspective of LLMs’ internal mechanism. Specifically, we visualize the feature representations of harmful prompts and boundary-safe prompts, where the latter are harmless but frequently rejected by LLMs. We find that existing defense methods fail to learn a precise boundary that distinguishes the feature representations of harmful and boundary-safe prompts, as shown in Fig. 1. In this way, boundary-safe representations close to harmful ones are inevitably affected during fine-tuning with these defense methods. Consequently, these boundary-safe representations are mistakenly treated as harmful, leading to the rejection of the corresponding prompts by LLMs.
@@
To reconcile the trade-off between defense robustness and usability, we propose X-Boundary that explicitly formulates the boundary between harmful and safe representations. Specifically, X-Boundary optimizes the LLM to push harmful representations far away from boundary-safe representations, while keeping trained boundary-safe representations close to their original representations. In this way, X-Boundary obtains a precise distinction boundary, and these harmful representations are further erased. Experimental results show that X-Boundary relatively reduces the attack success rate (ASR) of ten jailbreak attacks by over 70%, while lowering the ORR by approximately 20% compared to other defense methods, with almost no decline in general capability. Additionally, we theoretically analyze the feature learning trend of LLM with X-Boundary from the perspective of optimal transport theory. Theoretical analysis and experimental results indicate that X-Boundary achieves 22% improvement in the learning speed.
Recent studies safechain; R1_assessment suggest that large reasoning models (LRMs) with strong reasoning abilities and extended thinking processes may pose greater potential harm. To address this, we adapt both existing defense methods and X-Boundary to DeepSeek-R1 distilled reasoning models. On LRMs, existing methods either fail to establish effective defenses or severely impair the model’s reasoning capabilities. In contrast, X-Boundary outperforms other methods in defense effectiveness, while maintaining the average ORR below 10% and preserving 99% of reasoning ability. With its strong adaptability, we hope that X-Boundary can complement existing alignment methods to provide a more efficient and fine-grained defense, ultimately enhancing the prospects of deploying robust AI systems in diverse real-world applications.
2 The Trade-Off Between Defense Robustness and LLM Usability
We adapt and comprehensively evaluate four classic defense methods, i.e., Supervised Fine-Tuning (SFT) decoupled_sft; actor_attack, Direct Preference Optimization (DPO) dpo; red_queen, Gradient Ascent (GA) safe_unlearning; eraser, and Circuit Breaking (CB) circuit_breaker on Qwen2.5-7B-Instruct qwen2.5. To establish defense against single-turn and multi-turn attacks, we construct a mixed training dataset comprising single-turn data from circuit_breaker and multi-turn data curated from SafeMTData actor_attack. We evaluate the defense robustness of the four methods against single-turn attack harmbench and multi-turn attack actor_attack, as well as their impact on usability, i.e., over-refusal oktest and the decline of general capability human_eval. The evaluation metrics are the Attack Success Rate (ASR), Over-Refusal Rate (ORR), and Accuracy, respectively. A lower ASR indicates greater defense robustness against jailbreak attacks. Details on data construction, training settings, and evaluations are illustrated in Appendix D.1, Appendix D.2, and Appendix D.5, respectively.
Figure 2: The trade-off between defense robustness and LLM usability on Qwen2.5-7B-Instruct. The green points are trained with only single-turn defense data. The red points are trained with single-turn and multi-turn defense data.
@@ 2 The Trade-Off Between Defense Robustness and LLM Usability
Existing defense methods are suffering from a trade-off, where defense robustness improves while LLM usability declines. Fig 2 shows that existing methods can effectively reduce the ASR of jailbreak attacks after training with the aforementioned data. However, SFT, DPO, and GA even tend to severely compromise general capabilities when achieving good performance, commonly referred to as the “alignment tax” alignment_tax. For instance, SFT results in about 5% decrease in coding abilities. Moreover, all of these methods lead to severe over-refusal problems. In particular, the average ORR increases to more than 50% after GA. The high ORR reflects that these methods cannot precisely distinguish harmful queries and build effective defense mechanisms for them. Instead, they simply reduce the ASR by indiscriminately rejecting input queries, which is not trustworthy and undermines the model’s usability in real-world scenarios. Therefore, it is necessary to analyze the cause of usability decline and propose a more precise defense method to mitigate it while preserving robustness against jailbreaks.
Multi-Turn defense significantly exacerbates the trade-off. Fig. 2 shows that multi-turn attacks achieve higher ASR than single-turn attacks on the vanilla model, indicating that multi-turn defense is particularly challenging scaleAI_multi; crescendo. After incorporating multi-turn defense data into the training set, the data points in Fig. 2 overall shift towards the upper right, illustrating the increased difficulty in balancing defense robustness and LLM usability in multi-turn scenarios. Notably, the average ORR increases by 25.65% following multi-turn GA, while the decline in coding capability grows by 2.24% after multi-turn SFT. These findings highlight that the trade-off issue, especially in multi-turn scenarios, cannot be overlooked and demands urgent resolution.
Existing over-refusal mitigation methods fail to resolve the trade-off. To further explore the trade-off issue, we implement three existing over-refusal mitigation methods: System Prompt (SP)  oktest, Self-CD oktest, and Vector Ablation (VA) wang2024surgical. As shown in Table 4 in Appendix C.1, their effectiveness in reducing ORR is not noticeable in models fine-tuned with defense methods, and they substantially compromise defense robustness. Specifically, SP, Self-CD, and VA lead to increases of 20%, 7.5%, and 22.5% in multi-turn ASR, respectively, highlighting that they cannot reconcile the trade-off between minimizing ASR and maintaining usability.
3 X-Boundary: Optimize Exact Boundary to Balance Robustness and Usability
@@
In this section, we propose X-Boundary to mitigate the trade-off between defense robustness and LLM usability by explicitly formulating the distinction boundary. Section 3.1 analyzes the essential mechanism of decline in usability. Section 3.2 introduces the optimization objective of X-Boundary. Section 3.3 theoretically proves that X-Boundary may ease the learning difficulty and contribute to fast learning.
@@
Analysis of safety-usability trade-off from the perspective of interpretability mechanism. Existing defense methods circuit_breaker; zou2023representation typically improve the adversarial robustness of LLMs by intervening harmful feature representations $\{\mathcal{R}_{\mathcal{M}}\left(x_{i}^{h}\right)\}_{i=1}^{N}$. Specifically, SFT decoupled_sft and CB circuit_breaker remap harmful representations to refusal representations $\mathcal{R}_{\mathcal{M}}\left(x_{i}^{r}\right)$. In this process, these methods implicitly train LLMs to learn a boundary that distinguishes harmful representations and safe representations $\{\mathcal{R}_{\mathcal{M}}\left(x_{i}^{s}\right)\}_{i=1}^{N}$. However, Fig. 3 shows that the boundary learned through this implicit training is imprecise, with some boundary-safe representations $\{\mathcal{R}_{\mathcal{M}}\left(x_{i}^{b}\right)\}_{i=1}^{N}$ mixed with harmful representations rather than being clearly distinguished. In this way, these boundary-safe representations are mistakenly treated as harmful ones, leading LLMs to refuse the corresponding boundary-safe queries and ultimately reducing usability.
@@
3.2 Explicit Formulation for Distinction Representation Boundary
We propose X-Boundary to explicitly formulate the distinction boundary between safe and harmful representations. The key idea is to push harmful representations far away from boundary-safe representations through an explicit loss function, such that harmful representations can be effectively and precisely erased without disrupting safe ones. In this way, a balance between defense robustness and LLM usability can be achieved.
@@ 3.2 Explicit Formulation for Distinction Representation Boundary
To explicit formulate a precise distinction boundary, we propose separate loss $\mathcal{L}_{\texttt{s}}$ to increase the distance $\mathcal{D}$ between harmful representations $\{\mathcal{R}_{\mathcal{M}_{\theta}}\left(x_{i}^{h}\right)\}_{i=1}^{N}$ and boundary-safe representations $\{\mathcal{R}_{\mathcal{M}_{\texttt{ref}}}\left(x_{i}^{b}\right)\}_{i=1}^{N}$. Since most $\{\mathcal{R}_{\mathcal{M}_{\theta}}\left(x_{i}^{h}\right)\}_{i=1}^{N}$ will be remapped to $\{\mathcal{R}_{\mathcal{M}_{\theta}}\left(x_{i}^{r}\right)\}_{i=1}^{N}$ due to the following erasure operation, we can separate them by directly optimizing $\mathcal{R}_{\mathcal{M}_{\theta}}\left(x_{i}^{r}\right)$ to be orthogonal to $\mathcal{R}_{\mathcal{M}_{\texttt{ref}}}\left(x_{i}^{b}\right)$ as shown in Fig. 4:
$\mathcal{L}_{\texttt{s}}=\frac{1}{\left|D_{s}\right|}\sum_{i=1}^{\left|D_{s}\right|}\texttt{ReLU}\left(\texttt{cos}\left(\mathcal{R}_{\mathcal{M}_{\theta}}\left(x_{i}^{r}\right),\mathcal{R}_{\mathcal{M}_{\texttt{ref}}}\left(x_{i}^{b}\right)\right)\right)$ (1)
where $\mathcal{M}_{\theta}$ and $\mathcal{M}_{\texttt{ref}}$ denote the model under training and the reference model before training.
To establish robust defense against multi-turn attacks, we utilize erase loss $L_{e}$ to erase the representations of harmful QA pairs in $D_{\texttt{e}}$. $L_{e}$ optimizes $\mathcal{R}_{\mathcal{M}_{\theta}}\left(x_{i}^{h}\right)$ to be orthogonal to their original representations $\mathcal{R}_{\mathcal{M}_{\texttt{ref}}}\left(x_{i}^{h}\right)$ following circuit_breaker:
@@
To preserve usability of LLMs, we use retain loss $\mathcal{L}_{\texttt{r}}$ to maintain safe representations of data points in $D_{\texttt{r}}$. $\mathcal{L}_{\texttt{r}}$ minimizes the $\ell_{2}$ distance between trained representations and their original representations:
$\mathcal{L}_{\text{r}}=\frac{1}{\left|D_{r}\right|}\sum_{i=1}^{\left|D_{r}\right|}\left\|\mathcal{R}_{\mathcal{M}_{\theta}}\left(x_{i}\right)-\mathcal{R}_{\mathcal{M}_{\texttt{ref}}}\left(x_{i}\right)\right\|_{2}$ (3)
where $x_{i}$ represents a sample in retain set ($x_{i}\in D_{r}$). Notably, to maintain the existing refusal mechanism of LLMs, refusal responses $x_{r}$ to harmful queries are added into $D_{\texttt{r}}$. Therefore, most $\{\mathcal{R}_{\mathcal{M}_{\theta}}\left(x_{h}\right)\}_{i=1}^{N}$ are finally optimized to refusal representations $\{\mathcal{R}_{\mathcal{M}_{\theta}}\left(x_{r}\right)\}\}_{i=1}^{N}$ under the joint effect of $\mathcal{L}_{\texttt{e}}$ and $\mathcal{L}_{\texttt{r}}$.
@@
where $c_{r}$, $c_{e}$ and $c_{s}$ are adaptive loss coefficients following circuit_breaker; adaptive_loss. With the above optimization objective, X-Boundary can perform fine-grained optimization in the representation space to reconcile the trade-off between defense robustness and the usability of LLMs. The overall optimization process of X-boundary is shown as Algorithm 1 in Appendix B.